# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v2

@@ 4.1 Nonnegative and low-rank matrix completion
The first experiment we consider is nonnegative matrix completion/factorization (Lee and Seung, 1999), whose goal is to recover missing entries of a partially observed matrix using nonnegative entries, and we impose the additional structure requirement that the recovered matrix should be low-rank. Given a data matrix $M\in{\rm I}\!{\rm R}^{m\times n}$, we let $\Omega$ denote the entries $(i,j)$ oberserved and $P_{\Omega}$ denote the associated projection operation such that $P_{\Omega}(X)_{i,j}$ outputs $X_{i,j}$ if $(i,j)\in\Omega$ and 00 otherwise. We consider Eq. 1.1 with $p(X)\equiv 0$ and
$f(X)=\frac{\lambda_{1}}{2}\min_{Y\in{\rm I}\!{\rm R}^{m\times n}_{+}}\|X-Y\|^{2}_{F},\quad g(X)=\lambda_{2}\|X\|_{*},\quad h(X)=\frac{1}{2}\|P_{\Omega}(X-M)\|^{2}_{F},$
@@
where $\lambda_{1},\lambda_{2}\geq 0$ are weights for the respective terms, $\|\cdot\|_{*}$ is the nuclear norm, $\|\cdot\|_{F}$ is the Frobenius norm, and ${\rm I}\!{\rm R}^{m\times n}_{+}$ is the nonnegative orthant of ${\rm I}\!{\rm R}^{m\times n}$.
@@
We follow Toh and Yun (2010) to generate $M\in{\rm I}\!{\rm R}^{m\times n}$ with a specific rank $r$ as the product of an $m\times r$ and an $r\times n$ matrix whose entries are all identically and independently distributed as standard Gaussian, and then randomly select $s$ entries in uniform random to form $\Omega$. We in particular use $m=n$ and consider $(n,s)\in\{(100,1000),(500,10000)\}$, $r\in\{10,30\}$, $\lambda_{1}=5$ and $\lambda_{2}=10$. We set $\beta=\infty$ for our method due to the absence of the $p$ term. Since $f$ is not strongly convex, our theoretical results indicate that we should use $\tau<2$. We thus run our method with $\tau\in\{1,1.1,1.2,1.3,1.4,1.5,1.6,1.7,1.8,1.9\}$, [^2] We have also experimented with $\tau\in(0,1)$, but those values tend to give worse performance. We therefore omit these results in our experiments. and compare with the proximal gradient method (PG) by treating $f+h$ together as the smooth term. In addition, for this problem, our method with $\tau=1$ recovers the Davis-Yin splitting (DYS). All algorithms are run with a cap of $30,000$ iterations. We can clearly see from the results shown in Table 1 that our method constantly outperforms DYS and PG.
@@ 4.2 Cardinality-constrained problems
Our next experiment considers a least-square problem with a penalized form of the cardinality constraint $\|x\|_{0}\leq k$ for some given positive integer $k$. As argued in (Gotoh et al., 2018), this constraint is equivalent to $\|x\|_{1}-\|x\|_{(k)}=0$, where $\|x\|_{(k)}$ is the Ky Fan $k$-norm that computes the sum of the top $k$ largest elements of the element-wise absolute value of $x$, and we follow Gotoh et al. (2018) to take $\|x\|_{1}-\|x\|_{(k)}$ as a penalty in the objective function instead of enforcing the hard constraint. We also follow a common convention in machine learning to add in a small Tikhonov regularization of the form $\|x\|_{2}^{2}$ to improve the problem condition. This leads to the following setting for Eq. 1.1.
$f(x)=\frac{\lambda_{1}}{2}\|x\|_{2}^{2},\quad g(x)=\lambda_{2}\|x\|_{1},\quad p(x)=-\lambda_{2}\|x\|_{(k)},\quad h(x)=\frac{1}{2}\|Ax-b\|^{2}_{2}.$ (4.1)
@@
Given the existence of the $p$ term in this experiment, PG is not applicable, but by again treating $(f+h)$ as the smooth term, we can use the proximal DC algorithm described in Eq. 3.6. Our experiment uses publicly available real-world data [^3] Downloaded from https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. listed in Table 2. In this experiment, we set $\lambda_{1}=0.01$, $\lambda_{2}=0.005$, and $k=\lfloor n/10\rfloor$. As Eq. 3.49 does not hold for any $\tau\geq 2$, for our method we use the same range of $\tau$ as the previous experiment. All algorithms are run with a cap of $100,000$ iterations. The results are shown in Table 3. We can see that with appropriately selected $\tau$, our method performs similarly to proximal DC on ijcnn1, and is much more efficient than it on all the other datasets.
Data $m$ $n$
colon-cancer 62 2000
duke 38 7129
ijcnn1 49900 22
phishing 11055 68
heart 270 13
Table 2: Datasets used in the experiment of least-square problems.
@@
Table 3: Comparison between proximal DC (denoted by PDC) and our method for Eq. 4.1 * denotes that the maximum iteration of $100000$ is reached. $\tau=t$ denotes our method with the specified $\tau$.
Method colon duke ijcnn1 phishing heart
iter time iter time iter time iter time iter time
PDC 59796 63.88 19108 33.09 1367 2.97 *100000 *104.20 *100000 *27.56
$\tau=1$ 49934 53.09 13600 26.41 1502 3.33 *100000 *104.31 *100000 *24.96
$\tau=1.1$ 37384 39.76 8447 17.08 1650 3.69 *100000 *104.94 *100000 *25.31
$\tau=1.2$ 20352 21.71 5538 10.30 1834 4.06 *100000 *103.98 *100000 *25.70
$\tau=1.3$ 4772 5.07 5500 10.06 2066 4.60 *100000 *104.64 *100000 *26.02
$\tau=1.4$ 4672 4.97 5847 11.64 2370 5.26 *100000 *104.08 *100000 *31.02
$\tau=1.5$ 5148 5.48 6450 13.61 2787 6.19 *100000 *114.73 *100000 *34.09
$\tau=1.6$ 5968 6.39 7365 13.56 3396 7.55 28347 29.39 *100000 *30.16
$\tau=1.7$ 7335 7.84 8790 17.05 4380 9.71 32633 33.78 *100000 *31.07
$\tau=1.8$ 9881 10.52 11220 21.21 6257 13.92 42635 44.08 *100000 *33.36
$\tau=1.9$ 16308 17.54 16169 31.22 11457 25.25 68361 70.27 52222 14.46
5 Conclusion
@@
In this work, we presented a new splitting algorithm for the four-term optimization problem (1.1). Our algorithm generalizes the Davis-Yin splitting algorithm for three-term optimization to allow for an additional nonsmooth term. We derived stepsize estimates for the proposed algorithm that ensure global subsequential convergence to stationary points of the objective function. A notable implication of our results is the significant improvement in the upper bound estimates for the stepsize of the DYS algorithm that guarantee subsequential convergence. Our experiments demonstrated the applicability and effectiveness of our proposed algorithm.

### v4

@@ 4.1 Nonnegative and low-rank matrix completion
The first experiment we consider is nonnegative matrix completion/factorization (Lee and Seung, 1999), whose goal is to recover missing entries of a partially observed matrix using nonnegative entries, and we impose the additional structure requirement that the recovered matrix should be low-rank. Given a data matrix $M\in{\rm I}\!{\rm R}^{m\times n}$, we let $\Omega$ denote the entries $(i,j)$ observed and $P_{\Omega}$ denote the associated projection operation such that $P_{\Omega}(X)_{i,j}$ outputs $X_{i,j}$ if $(i,j)\in\Omega$ and 00 otherwise. We consider Eq. 1.1 with $p(X)\equiv 0$ and
$f(X)=\frac{\lambda_{1}}{2}\min_{Y\in{\rm I}\!{\rm R}^{m\times n}_{+}}\|X-Y\|^{2}_{F},\quad g(X)=\lambda_{2}\|X\|_{*},\quad h(X)=\frac{1}{2}\|P_{\Omega}(X-M)\|^{2}_{F},$ (4.2)
@@
where $\lambda_{1},\lambda_{2}\geq 0$ are weights for the respective terms, $\|\cdot\|_{*}$ is the nuclear norm, $\|\cdot\|_{F}$ is the Frobenius norm, and ${\rm I}\!{\rm R}^{m\times n}_{+}$ is the nonnegative orthant of ${\rm I}\!{\rm R}^{m\times n}$. We can easily see that Eq. 4.2 satisfies our assumptions with $L_{f}=\lambda_{1},L_{h}=1$, and $\rho_{p}=\rho_{f}=\rho_{h}=\sigma_{h}=0$.
@@
We follow Toh and Yun (2010) to generate $M\in{\rm I}\!{\rm R}^{m\times n}$ with a specific rank $r$ as the product of an $m\times r$ and an $r\times n$ matrix whose entries are all identically and independently distributed as standard Gaussian, and then randomly select $s$ entries in uniform random to form $\Omega$. We in particular use $m=n$ and consider $(n,s)\in\{(100,1000),(300,10000)\}$, $r\in\{10,30\}$, $\lambda_{1}=10$ and $\lambda_{2}=5$. We set $\beta=\infty$ for all applicable methods due to the absence of the $p$ term. All algorithms are run with a cap of $30,000$ iterations.
4.1.1 Comparison of different bounds for $\alpha$
We first use our method with $\tau=1$ to recover the DYS algorithm, and compare our obtained bound for $\bar{\alpha}$ with those in Bian and Zhang (2021); Dao et al. (2024) for the same setting. We use the same factor of $0.99$ for all the obtained upper bounds to ensure that strict inequality holds, which is essential to guarantee theoretical convergence. The result in Table 1 shows that a larger $\alpha$ leads to faster convergence, and therefore our new upper bound for $\alpha$ provides not only theoretical but also practical improvements.
Table 1: Comparison between different bounds for $\alpha$ in DYS using nonnegative and low-rank matrix completion. * denotes that the maximum iteration of $30000$ is reached. The respective values of $\bar{\alpha}$ in this experiment are: Bian and Zhang (2021): $0.0212$, Dao et al. (2024): $0.0645$, ours: $0.09$.
Bound $n=100,r=10$ $n=100,r=30$ $n=300,r=10$ $n=300,r=30$
iter time iter time iter time iter time
Bian and Zhang (2021) 14371 17.37 27335 32.64 14937 154.11 *30000 *322.91
Dao et al. (2024) 9087 11.89 17283 21.70 9445 100.31 21929 229.87
Ours 6516 8.71 12394 16.47 6773 73.02 15725 167.23
4.1.2 Comparison with other methods
We next compare our method with different $\tau$’s. Since $f$ is not strongly convex, our theoretical results indicate that we should use $\tau<2$. We thus run our method with $\tau\in\{1,1.1,\cdots,1.9\}.$ [^7] We have also experimented with $\tau\in\{0.5,0.6,0.7,0.8,0.9\}$, but those values give worse performance, with smaller $\tau$ leading to slower convergence. We therefore omit these results in our experiments.
We compare our method with the proximal gradient method (PG) by treating $f+h$ together as the smooth term, and we set the stepsize to be $0.99/(L_{{}_{\rm{f}}}+L_{{}_{\rm{h}}})$ following standard practice. In addition, as noticed in the previous part, for this problem, our method with $\tau=1$ recovers the Davis-Yin splitting (DYS). According to the result of the previous experiment, for DYS we use the new upper bound for $\alpha$ obtained in this work. We can clearly see from the results in Table 2 that our method with $\tau\in(1,1.9)$ constantly outperforms DYS (with our improved upper bound for $\alpha$) and PG, and DYS and PG are performing very similarly in this experiment.
@@ 4.2 Cardinality-constrained problems
Our next experiment considers a least-squares problem with a penalized form of the cardinality constraint $\|x\|_{0}\leq k$ for some given positive integer $k$. As argued in (Gotoh et al., 2018), this constraint is equivalent to $\|x\|_{1}-\|x\|_{(k)}=0$, where $\|x\|_{(k)}$ is the Ky Fan $k$-norm that computes the sum of the top $k$ largest elements of the element-wise absolute value of $x$, and we follow Gotoh et al. (2018) to take $\|x\|_{1}-\|x\|_{(k)}$ as a penalty in the objective function instead of enforcing the hard constraint. We also follow a common convention in machine learning to introduce a Tikhonov regularization of the form $\|x\|_{2}^{2}$ to improve the problem condition. This leads to the following setting for Eq. 1.1.
$f(x)=\frac{1}{2}\|Ax-b\|^{2}_{2},\quad g(x)=\lambda_{2}\|x\|_{1},\quad p(x)=-\lambda_{2}\|x\|_{(k)},\quad h(x)=\frac{\lambda_{1}}{2}\|x\|_{2}^{2}.$ (4.3)
For Eq. 4.3, we have the following parameter estimates: $L_{f}=\|A^{\top}A\|_{2},L_{h}=\lambda_{1},\rho_{f}=\rho_{h}=\rho_{p}=0$, and $\sigma_{f}$ equals to the smallest eigenvalue of $A^{\top}A$.
@@
Given the existence of the $p$ term in this experiment, PG is not applicable, but by again treating $(f+h)$ as the smooth term, we can use the proximal DC algorithm described in Eq. 3.6. In this experiment, we set $\lambda_{1}=5$, $\lambda_{2}=10$, and $k=\lfloor n/10\rfloor$. As Eq. 3.41 does not hold for any $\tau\geq 2$, for our method we use the same range of $\tau$ as the previous experiment. All algorithms are run with a cap of $100,000$ iterations. Results on publicly available real-world data sets are shown in Table 3.888Downloaded from https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/.
In terms of the number of iterations, we can see that our method with $\tau>1$ is more efficient than PDC, and before reaching $\tau=2$, the number of iterations of our method is roughly proportional to $1/\tau$. For those datasets on which $\tau=2$ is feasible, we see that the per-iteration decrease for $\tau=2$ calculated in the proof of Theorem 3.12 is very close to 00, which already suggested a slow convergence. We therefore see that the value $c(\alpha)$ could be a good indicator for the speed of practical convergence as well.
On the other hand, in terms of the total running time, we observe that by utilizing the problem structure wisely, our method also achieves significantly lower time cost per iteration. This shows that although it is always possible to solve the problem Eq. 1.1 by existing methods by considering less problem structures (in the worst case, one can always just consider it as a nonsmooth optimization problem and use a subgradient method to solve it), wisely exploiting problem structures can lead to much higher practical efficiency.
@@
Table 3: Comparison between proximal DC (denoted by PDC) and our method for Eq. 4.3. * denotes that the maximum iteration of $100000$ is reached, $\tau=t$ denotes our method with the specified $\tau$, and -- denotes that this value of $\tau$ is not applicable for this problem.
a9a cpusmall_scale ijcnn1 phishing heart
$m=$ $n=$ $m=$ $n=$ $m=$ $n=$ $m=$ $n=$ $m=$ $n=$
32561 123 8192 12 49900 22 11055 68 270 13
Method iter time iter time iter time iter time iter time
PDC 91859 67.21 37132 5.64 2942 1.92 7283 3.27 *100000 *2.42
$\tau=1$ 91865 6.80 37133 1.34 2948 0.16 7282 0.40 *100000 *1.77
$\tau=1.1$ 83514 6.02 33757 1.23 2680 0.14 6620 0.36 91675 1.56
$\tau=1.2$ 76555 5.74 30944 1.18 2456 0.14 6069 0.34 84035 1.48
$\tau=1.3$ 70664 5.05 28564 1.03 2267 0.13 5602 0.32 77571 1.42
$\tau=1.4$ 65618 4.85 26524 1.10 2105 0.13 5202 0.30 72030 1.36
$\tau=1.5$ 61242 4.64 24756 0.97 1965 0.12 4856 0.28 67228 1.20
$\tau=1.6$ 57415 4.13 23209 0.86 1842 0.11 4553 0.26 63027 1.13
$\tau=1.7$ 54038 4.09 21844 0.76 1733 0.11 4285 0.24 59319 1.17
$\tau=1.8$ 51037 3.88 20630 0.74 1637 0.10 4047 0.23 56024 1.01
$\tau=1.9$ 48350 3.63 19545 0.68 1550 0.09 3835 0.22 53075 0.94
$\tau=2$ 1746 0.11 - - 97789 1.83
5 Conclusion and Future Work
@@
In this work, we presented a new splitting algorithm for the four-term optimization problem (1.1). Our algorithm generalizes the Davis-Yin splitting algorithm for three-term optimization to allow for an additional nonsmooth term. We derived stepsize estimates for the proposed algorithm that ensure global subsequential convergence to stationary points of the objective function. A notable implication of our results is the significant improvement in the upper bound estimates for the stepsize of the DYS algorithm that guarantee subsequential convergence.
Our numerical experiments demonstrate that larger stepsizes can significantly accelerate the convergence of our proposed algorithm, emphasizing the critical role of stepsize selection in practical performance. Previous works, such as Wright et al. (2009); Lu and Zhou (2019), have shown that incorporating adaptive stepsizes safeguarded by line search techniques can improve the efficiency of PG and PDC, which are special cases of our framework. Our preliminary experiments suggest that such techniques, while not preserving iteration complexity guarantees, can substantially enhance empirical convergence speed. Motivated by these findings, we are actively exploring extensions of adaptive stepsize strategies to our more general algorithm, aiming to further enhance efficiency while maintaining solution quality and the convergence rate guarantees.